Solve the system using substitution or elimination:
2x-4=-2
4x+8y=0

Answer by gems(12)

2x-4=-2
4x+8y=0
using substitution 2x-4=-2
2x=-2+4
2x=2
x=1
if x=1
using 2nd equation 4x+8y=0
4(1)+8y=0
4+8y=0
8y=-4
y=-4/8
y=-1/2